# Cogenesis of visible and dark matter in a scotogenic model

### v3

@@ 1 Introduction
The origin of matter in the Universe remains a mystery within the standard paradigm of particle physics and cosmology. However, we have robust evidences to believe in the existence of two components of the gravitating matter viz. the luminous baryons that make up the stars that light up the night sky ($15\%$) and the dark matter (DM) that form the galactic halos ($85\%$) that hosts these stars Cirelli:2024ssz. This is further compounded with the observations of matter antimatter asymmetry at the cosmological scales as evidenced by the CMB/BBN studies Planck:2018vyg; ParticleDataGroup:2024cfk. Today we have a sound theoretical and ever increasing experimental evidence of the zoo of particles that populate the visible sector, however the particulate nature of the dark sector remains a speculation. It is possible to consider scenario where the dark sector is dominated by primordial black holes Carr:2020xqk, MACHOs Griest:1990vu; Gates:1995js or may reflects a modified description of gravity Milgrom:1983ca. In this work we will presume a multipartite dark sector in similar spirit to the visible particle spectrum. A common origin of dark matter and visible matter has the mandate to explain both the origin of the relative density of the two sectors $\Omega_{DM}/\Omega_{b}\sim 5$ and the observed matter-antimatter asymmetry in the visible Universe $\eta_{b}\sim 10^{-10}$ within a natural unified theory.
@@
Cogenesis is a proposed framework that addresses these issues by linking the process of baryogenesis with dark matter production in the early Universe Davoudiasl:2012uw; Dasgupta:2016odo; Bernal:2017zvx; Borah:2018uci; Mahanta:2019gfe; Dasgupta:2019lha; Cui:2020dly; Chu:2021qwk. In this article, we provide a novel version of cogenesis originating from the scotogenic neutrino mass model giving rise to a multipartite dark sector. In contrast to the usual approach to consider an asymmetric freeze-in of baryons and thermal freeze-out of DM Kashiwase:2012xd; Hugle:2018qbw; Baumholzer:2018sfb; Heisig:2024mwr, we present a mechanism where the freeze-in production of visible and a non thermal dark sector are conjoint to the same CP violating decay of a $\mathbb{Z}_{2}$-odd heavy sterile neutrino within the minimal scotogenic framework. The CP violation of the decay dictates the present day baryon asymmetry through leptogenesis. The dark matter is generated by subsequent decay of the dark sector particle to the lightest stable state as sketched in Fig. 1. This decay is delayed due to the competition between the scattering and decay processes that is crucial in determining the relic density for the freeze-in dark matter to remain consistent with cosmological constraints from large scale structure formation.
@@
We sketch the minimal version of this framework within a scotogenic model having three $\mathbb{Z}_{2}$-odd right handed neutrinos (RHNs) and a single $\mathbb{Z}_{2}$-odd scalar doublet. With the masses of the heavier states set around $\mathcal{O}(10^{9})$ GeV, we demonstrate how this framework can simultaneously generate the observed baryon asymmetry and a freeze-in dark matter in the sub-GeV mass scale that saturates the relic density while remaining consistent with the neutrino oscillation parameters Esteban:2020cvm.
@@ 2 Model
In this section we briefly review the minimal Scotogenic model Ma:2006km. We make minimal extension of the SM by including a $\mathbb{Z}_{2}$-odd sector composed of a scalar $SU(2)_{L}$ doublet $(\eta)$ and a set of hierarchical heavy sterile neutrinos $(N_{i}).$ The charges of all the relevant particles are summarised in Table 1 and we will assume $M_{N_{3}}>M_{N_{2}}\gg M_{N_{1}}$. The lepton number violating Yukawa interactions consistent with the charge assignments of Table 1 can be written as
@@
where in contrast to the usual SM Higgs sector, the positive mass term of the inert scalar doublet preserves the $\mathbb{Z}_{2}$ symmetry by preventing the inert doublet from obtaining a vacuum expectation value (vev). The $\mathbb{Z}_{2}$-odd state $\eta$ couples to the visible sector through the Higgs portal couplings $\lambda_{3/4/5}.$ In the scenario where $\eta$ is the lightest dark sector state we can map this to a viable WIMP like Higgs portal DM and has been studied extensively in the literature Banerjee:2021anv; Banerjee:2021hal. However in this article we will consider it to be the next-to-lightest stable particle (NLSP) in the dark sector. Within this context we will demonstrate that the Higgs portal couplings play a crucial role in thermalising the NLSP before it decays to the DM state. These interaction of the NLSP with the primordial soup plays a key role in producing a viable sub-GeV dark matter state. Explicitly in terms of the components the scalar fields can be written as
@@
The masses of the SM neutrinos are generated radiatively at one loop as shown in Fig. 2 and is given by Escribano:2020iqq
@@
which is directly proportional to the non-self hermitian part of the potential given in Eq. 2 as stated above. The solar and atmospheric neutrino mass scales are generated from the heavy seesaw masses of $N_{2}$ and $N_{3}$. The Yukawa structure in Eq. 1 has 18 parameters and are responsible to generate the neutrino oscillation parameters Esteban:2020cvm. The lightest RHN $N_{1}$ which emerges as the DM candidate have a negligible contribution towards the neutrino mass due to its tiny Yukawa couplings Molinaro:2014lfa and will be demonstrated to be produced by late time decay of the $\eta$ state producing a cosmologically viable sub-GeV freeze-in dark matter.
@@ 3 Cogenesis
Figure 3: Tree level (3(a)) and $N_{3}$ the leading one loop (3(b) and 3(c)) decay of $N_{2}$.
@@
where $x=M_{N_{3}}^{2}/M_{N_{2}}^{2}$. Note that the heaviest RHN ($N_{3}$) appearing in the loop is crucial for the generation of a sizable primordial asymmetry as the contribution coming from the lightest RHN ($N_{1}$) is numerically insignificant. Once these asymmetries are generated they evolve according to the collisional Boltzmann equations to produce a remnant present day lepton asymmetry eventually translating to a net baryon asymmetry via standard sphaleron processes Khlebnikov:1988sr; Harvey:1990qw.
@@
$5\times 10^{10}$ $5\times 10^{9}$ $5\times 10^{8}$ 8.114 0 $3.046\times 10^{-2}$
@@
where the ratio of Bessel functions accounts for the dilution factor Buchmuller:2004nz arising from the expansion of the Universe. The reaction densities for generic $2\leftrightarrows 2$ scattering processes in the center of mass frame is denoted as $\gamma^{X}_{Y}$ and is given by
@@
where $\widehat{\sigma}$ is the reduced cross section Luty:1992un.
@@
$33.535\degree$ $50.314\degree$ $8.918\degree$
@@
$1.072\times 10^{-41}$ $3.160\times 10^{-50}$ $1.439\times 10^{-55}$ cm
@@
We choose a benchmark point given in Table 3 to illustrate the dynamics of cogenesis in the minimal scotogenic model. The decay reaction densities of the inert scalar doublet as well as its scatterings with the Higgs are shown in Fig. 4(a). In presence of larger quartic couplings and relatively smaller decay width of $\eta$, the scattering processes tend to keep inert doublet in thermal equilibrium delaying its decay into a lepton and $N_{1}$. The equality of these reaction densities occurs at $z\sim 250$ denoted by a vertical dashed gray line in 4(a) after which the decay reaction density takes over populating the dark sector with $N_{1}$. All other reaction densities are shown in Fig. 4(b) including the $\Delta L=2$ scattering processes which modules the asymmetry between the dark sector and the visible sector. The evolution of various comoving number densities are shown in Fig. 5 with respect to the dimensionless variable $z$. As can be seen from the figure the CP violating decay of $N_{2}$ generates an asymmetry in the lepton sector and in the inert doublet. The lepton asymmetry gets converted into the baryon asymmetry via standard sphaleron processes yielding the observed BAU measured by Planck Planck:2018vyg shown via the narrow horizontal shaded band with the numerical value given in Table 4. However, the produced inert doublet rapidly thermalises with the primordial plasma owing to its high scattering rate regulated by its couplings with the SM Higgs $(\lambda_{3})$ delaying its decay. As the Universe expands further, the flux suppression increases and around $z\sim 250$ the decay process starts to dominate over the scattering processes populating the dark sector with stable DM candidate $N_{1}$. The $\Delta L=2$ scattering processes modulates the asymmetry between the dark sector and the visible sector.
@@
Figure 6: Parameter space yielding correct BAU and relic abundance while satisfying the structure formation bounds. The benchmark point given in Table 3 is indicated by ✖.
@@
In our framework, the main constraints on the minimum mass of DM particles arise from the cosmic microwave background (CMB) and big bang nucleosynthesis (BBN) observations. For DM that annihilates through s-wave processes, energy injection from annihilation can alter the CMB spectrum, leading to a lower bound on the DM mass, typically $m_{\chi}\gtrsim\mathcal{O}(10)\,{\rm GeV}$ Leane:2018kjk. For other scenarios, such as p-wave annihilation, BBN provides the strongest constraint since relativistic DM during BBN can alter the primordial element abundances, requiring the DM to be heavier than the MeV scale Sabti:2019mhn.
@@
In our framework, DM is produced via the freeze-in mechanism, where the coupling to the Standard Model is so feeble that it does not affect the CMB. Additionally, since DM is never in thermal equilibrium with the primordial plasma, its abundance does not impact BBN either. Therefore, for freeze-in DM, the lower mass limit mainly comes from structure formation bounds. Observations of dwarf galaxies, which are kpc scales, imply that the DM de-Broglie wavelength must be smaller than these structures, giving $m_{\chi}\gtrsim 10^{-21}\,{\rm eV}$ Zimmermann:2024xvd. If the DM is fermionic, the Pauli exclusion principle comes into play which pushes the lower bound at keV scales Tremaine:1979we; Alvey:2020xsk. These limits are independent of the DM production mechanism and depend mainly on whether the DM is bosonic or fermionic. In the typical freeze-in scenarios, DM is mainly produced when the temperature of the universe is around the mass scale of the mother particle. After production, DM free-streams, and if it is too light, its kinetic energy can wash out structures. As calculated in DEramo:2020gpr, using phase space distributions and Lyman-$\alpha$ forest data, the lower mass bound for such DM is found to be larger than $\mathcal{O}(1)$ keV.
@@
In our framework, the production of the dark matter particle $(N_{1})$ differs slightly from the usual freeze-in scenario. Both $N_{2}$ and the scalar $\eta$ contribute to the production of $N_{1}$ during cosmic evolution, with the dominant contribution arising from the decay of the complex scalar $\eta$. Importantly, $\eta$ does not decay immediately after decoupling from the plasma; instead, most of the $N_{1}$ population is generated through its late decay. As shown in Decant:2021mhj, this leads to a momentum distribution of $N_{1}$ that deviates from the standard freeze-in case. Since the momentum distribution governs the free-streaming length of DM and thus impacts structure formation, the corresponding constraints are modified by the decay rate of the parent particle and the bound is given by Decant:2021mhj
@@
where the dimensionless parameter $R_{\eta}=(M_{\rm Pl,red}\,\Gamma_{\eta})/M_{\eta}^{2}$, with $M_{\rm Pl,red}$ denoting the reduced Planck mass, $\Gamma_{\eta}$ is the decay width of $\eta$ and $g_{*S}$ is the effective entropy degrees of freedom at the relevant decay temperature ($T_{\text{decay}}$). In Eq. 13, $m_{\rm lim}$ corresponds to the keV-scale lower bounds from Lyman-$\alpha$ and $\Delta N_{\rm eff}$ probes, as listed in Table 1 of Decant:2021mhj. We adopt the most stringent value, $m_{\rm lim}=3.9$ keV. With a typical scattering cross-section with SM leptons below $\mathcal{O}(10^{-50})\,\text{cm}^{2}$ they easily remain below the present and future sensitivity of direct detection experiments XENON:2019gfn.
In our framework, the dark matter is the lightest RHN ($N_{1}$) produced from the delayed decay of the inert doublet. The mass of DM is set by considering that it accounts for the entire relic abundance of DM i.e $\Omega_{\text{DM}}h^{2}=0.12$ Planck:2018vyg and is found to be $M_{N_{1}}\simeq 9.6$ MeV which is larger than the lower mass bound on FIMP mass which is calculated to be $\simeq 4.3$ MeV according to Eq. 13 as explicitly mentioned in Table 4. To explore the parameter region for successful cogenesis while remaining consistent with the structure formation bounds, we perform an extensive numerical scan utilizing a dedicated Markhov Chain Monte Carlo algorithm following Sarazin:2021nwo over the 18 free Yukawa couplings and two quartic couplings as given in Table 2. We fix the mass scales as $M_{3}=5\times 10^{10}$ GeV, $M_{2}=5\times 10^{9}$, $M_{\eta}=5\times 10^{8}$ GeV, quartic coupling $\lambda_{4}=0$ and the allowed parameter region is shown in Fig. 6 where the benchmark point is denoted with ✖. The present and forthcoming collider experiments are not expected to provide a direct probe for the heavy states within this framework.

### v4

@@ 1 Introduction
The origin of matter in the Universe remains a mystery within the standard paradigm of particle physics and cosmology. However, we have robust evidences to believe in the existence of two components of the gravitating matter viz. the luminous baryons that make up the stars that light up the night sky ($15\%$) and the dark matter (DM) that form the galactic halos ($85\%$) that hosts these stars Cirelli et al. (2024). This is further compounded with the observations of matter antimatter asymmetry at the cosmological scales as evidenced by the CMB/BBN studies Aghanim and others (2020); Navas and others (2024). Today we have a sound theoretical and ever increasing experimental evidence of the zoo of particles that populate the visible sector, however the particulate nature of the dark sector remains a speculation. It is possible to consider scenario where the dark sector is dominated by primordial black holes Carr and Kuhnel (2020), MACHOs Griest (1991); Gates et al. (1996) or may reflects a modified description of gravity Milgrom (1983). In this work we will presume a multipartite dark sector in similar spirit to the visible particle spectrum. A common origin of dark matter and visible matter has the mandate to explain both the origin of the relative density of the two sectors $\Omega_{DM}/\Omega_{b}\sim 5$ and the observed matter-antimatter asymmetry in the visible Universe $\eta_{b}\sim 10^{-10}$ within a natural unified theory.
@@
Cogenesis is a proposed framework that addresses these issues by linking the process of baryogenesis with dark matter production in the early Universe Davoudiasl and Mohapatra (2012); Dasgupta et al. (2016); Bernal and Fong (2017); Borah et al. (2020); Mahanta and Borah (2019); Dasgupta et al. (2020); Cui and Shamma (2020); Chu et al. (2022). In this article, we provide a novel version of cogenesis originating from the scotogenic neutrino mass model giving rise to a multipartite dark sector. In contrast to the usual approach to consider an asymmetric freeze-in of baryons and thermal freeze-out of DM Kashiwase and Suematsu (2012); Hugle et al. (2018); Baumholzer et al. (2018); Heisig (2024), we present a mechanism where the freeze-in production of visible and a non thermal dark sector are conjoint to the same CP violating decay of a $\mathbb{Z}_{2}$-odd heavy sterile neutrino within the minimal scotogenic framework. The CP violation of the decay dictates the present day baryon asymmetry through leptogenesis. The dark matter is generated by subsequent decay of the dark sector particle to the lightest stable state as sketched in Fig. 1. This decay is delayed due to the competition between the scattering and decay processes that is crucial in determining the relic density for the freeze-in dark matter to remain consistent with cosmological constraints from large scale structure formation.
@@
We sketch the minimal version of this framework within a scotogenic model having three $\mathbb{Z}_{2}$-odd right handed neutrinos (RHNs) and a single $\mathbb{Z}_{2}$-odd scalar doublet. With the masses of the heavier states set around $\mathcal{O}(10^{9})$ GeV, we demonstrate how this framework can simultaneously generate the observed baryon asymmetry and a freeze-in dark matter in the sub-GeV mass scale that saturates the relic density while remaining consistent with the neutrino oscillation parameters Esteban et al. (2020).
@@ 2 Model
In this section we briefly review the minimal Scotogenic model Ma (2006). We make minimal extension of the SM by including a $\mathbb{Z}_{2}$-odd sector composed of a scalar $SU(2)_{L}$ doublet $(\eta)$ and a set of hierarchical heavy sterile neutrinos $(N_{i}).$ The charges of all the relevant particles are summarised in Table 1 and we will assume $M_{N_{3}}>M_{N_{2}}\gg M_{N_{1}}$. The lepton number violating Yukawa interactions consistent with the charge assignments of Table 1 can be written as
@@
where in contrast to the usual SM Higgs sector, the positive mass term of the inert scalar doublet preserves the $\mathbb{Z}_{2}$ symmetry by preventing the inert doublet from obtaining a vacuum expectation value (vev). The $\mathbb{Z}_{2}$-odd state $\eta$ couples to the visible sector through the Higgs portal couplings $\lambda_{3/4/5}.$ In the scenario where $\eta$ is the lightest dark sector state we can map this to a viable weakly interacting massive particle-like Higgs portal DM and has been studied extensively in the literature Banerjee et al. (2021a, b). However in this article we will consider it to be the next-to-lightest stable particle (NLSP) in the dark sector. Within this context we will demonstrate that the Higgs portal couplings play a crucial role in thermalising the NLSP before it decays to the DM state. These interaction of the NLSP with the primordial soup plays a key role in producing a viable sub-GeV dark matter state. Explicitly in terms of the components the scalar fields can be written as
@@
The masses of the SM neutrinos are generated radiatively at one loop as shown in Fig. 2 and is given by Escribano et al. (2020)
@@
which is directly proportional to the non-self hermitian part of the potential given in Eq. 2 as stated above. The solar and atmospheric neutrino mass scales are generated from the heavy seesaw masses of $N_{2}$ and $N_{3}$. The Yukawa structure in Eq. 1 has 18 parameters and are responsible to generate the neutrino oscillation parameters Esteban et al. (2020). The lightest RHN $N_{1}$ which emerges as the DM candidate have a negligible contribution towards the neutrino mass due to its tiny Yukawa couplings Molinaro et al. (2014) and will be demonstrated to be produced by late time decay of the $\eta$ state producing a cosmologically viable sub-GeV freeze-in dark matter.
@@ 3 Cogenesis
Figure 3: Tree level (3(a)) and the leading one loop (3(b) and 3(c)) decay of $N_{2}$ involving $N_{3}$.
@@
where $x=M_{N_{3}}^{2}/M_{N_{2}}^{2}$. Note that the heaviest RHN ($N_{3}$) appearing in the loop is crucial for the generation of a sizable primordial asymmetry as the contribution coming from the lightest RHN ($N_{1}$) is numerically insignificant. Once these asymmetries are generated they evolve according to the collisional Boltzmann equations to produce a remnant present day lepton asymmetry eventually translating to a net baryon asymmetry via standard sphaleron processes Khlebnikov and Shaposhnikov (1988); Harvey and Turner (1990).
@@
$5\times 10^{10}$ $5\times 10^{9}$ $5\times 10^{8}$ 3.276 0 $3.096\times 10^{-2}$
@@
where the ratio of Bessel functions accounts for the dilution factor Buchmuller et al. (2005) arising from the expansion of the Universe. The reaction densities for generic $2\leftrightarrows 2$ scattering processes in the center of mass frame is denoted as $\gamma^{X}_{Y}$ and is given by
@@
where $\widehat{\sigma}$ is the reduced cross section Luty (1992).
@@
$31.333\degree$ $42.279\degree$ $8.343\degree$
@@
$1.157\times 10^{-41}$ $3.380\times 10^{-50}$ $1.572\times 10^{-55}$ cm
@@
We choose a benchmark point given in Table 3 to illustrate the dynamics of cogenesis in the minimal scotogenic model. The decay reaction densities of the inert scalar doublet as well as its scatterings with the Higgs are shown in Fig. 4(a). In presence of larger quartic couplings and relatively smaller decay width of $\eta$, the scattering processes tend to keep inert doublet in thermal equilibrium delaying its decay into a lepton and $N_{1}$. The equality of these reaction densities occurs at $z\sim 250$ denoted by a vertical dashed gray line in Fig. 4(a) after which the decay reaction density takes over populating the dark sector with $N_{1}$. All other reaction densities are shown in Fig. 4(b) including the $\Delta L=2$ scattering processes which modules the asymmetry between the dark sector and the visible sector. The evolution of various comoving number densities are shown in Fig. 5 with respect to the dimensionless variable $z$. As can be seen from the figure the CP violating decay of $N_{2}$ generates an asymmetry in the lepton sector and in the inert doublet. The lepton asymmetry gets converted into the baryon asymmetry via standard sphaleron processes yielding the observed BAU measured by Planck Aghanim and others (2020) shown via the narrow horizontal shaded band with the numerical value given in Table 4. However, the produced inert doublet rapidly thermalises with the primordial plasma owing to its high scattering rate regulated by its couplings with the SM Higgs $(\lambda_{3})$ delaying its decay. As the Universe expands further, the flux suppression increases and around $z\sim 250$ the decay process starts to dominate over the scattering processes populating the dark sector with stable DM candidate $N_{1}$. The $\Delta L=2$ scattering processes modulates the asymmetry between the dark sector and the visible sector.
@@
Figure 6: Parameter space yielding correct BAU and relic abundance while satisfying the structure formation bounds as presented in terms of the Yukawa couplings of (6(a)) $N_{1}$, (6(b)) $N_{2}$, (6(c)) $N_{3}$ RHN and (6(d)) quartic couplings between Higgs and the inert scalar doublet. The benchmark point in Table III is indicated by ✖.
@@
In our framework, the main constraints on the minimum mass of DM particles arise from the cosmic microwave background (CMB) and big bang nucleosynthesis (BBN) observations. For DM that annihilates through s-wave processes, energy injection from annihilation can alter the CMB spectrum, leading to a lower bound on the DM mass, typically $m_{\chi}\gtrsim\mathcal{O}(10)\,{\rm GeV}$ Leane et al. (2018). For other scenarios, such as p-wave annihilation, BBN provides the strongest constraint since relativistic DM during BBN can alter the primordial element abundances, requiring the DM to be heavier than the MeV scale Sabti et al. (2020).
@@
In our framework, DM is produced via the freeze-in mechanism, where the coupling to the Standard Model is so feeble that it does not affect the CMB. Additionally, since DM is never in thermal equilibrium with the primordial plasma, its abundance does not impact BBN either. Therefore, for freeze-in DM, the lower mass limit mainly comes from structure formation bounds. Observations of dwarf galaxies, which are kpc scales, imply that the DM de-Broglie wavelength must be smaller than these structures, giving $m_{\chi}\gtrsim 10^{-21}\,{\rm eV}$ Zimmermann et al. (2025). If the DM is fermionic, the Pauli exclusion principle comes into play which pushes the lower bound at keV scales Tremaine and Gunn (1979); Alvey et al. (2021). These limits are independent of the DM production mechanism and depend mainly on whether the DM is bosonic or fermionic. In the typical freeze-in scenarios, DM is mainly produced when the temperature of the universe is around the mass scale of the mother particle. After production, DM free-streams, and if it is too light, its kinetic energy can wash out structures. As calculated in D’Eramo and Lenoci (2021), using phase space distributions and Lyman-$\alpha$ forest data, the lower mass bound for such DM is found to be larger than $\mathcal{O}(1)$ keV.
@@
In our framework, the production of the dark matter particle $(N_{1})$ differs slightly from the usual freeze-in scenario. Both $N_{2}$ and the scalar $\eta$ contribute to the production of $N_{1}$ during cosmic evolution, with the dominant contribution arising from the decay of the complex scalar $\eta$. Importantly, $\eta$ does not decay immediately after decoupling from the plasma; instead, most of the $N_{1}$ population is generated through its late decay. As shown in Decant et al. (2022), this leads to a momentum distribution of $N_{1}$ that deviates from the standard freeze-in case. Since the momentum distribution governs the free-streaming length of DM and thus impacts structure formation, the corresponding constraints are modified by the decay rate of the parent particle and the bound is given by Decant et al. (2022)
@@
where the dimensionless parameter $R_{\eta}=(M_{\rm Pl,red}\,\Gamma_{\eta})/M_{\eta}^{2}$, with $M_{\rm Pl,red}$ denoting the reduced Planck mass, $\Gamma_{\eta}$ is the decay width of $\eta$ and $g_{*S}$ is the effective entropy degrees of freedom at the relevant decay temperature ($T_{\text{decay}}$). In Eq. 13, $m_{\rm lim}$ corresponds to the keV-scale lower bounds from Lyman-$\alpha$ and $\Delta N_{\rm eff}$ probes, as listed in Table 1 of Decant et al. (2022). We adopt the most stringent value, $m_{\rm lim}=3.9$ keV. With a typical scattering cross-section with SM leptons below $\mathcal{O}(10^{-50})\,\text{cm}^{2}$ they easily remain below the present and future sensitivity of direct detection experiments Aprile and others (2019).
In our framework, the dark matter is the lightest RHN ($N_{1}$) produced from the delayed decay of the inert doublet. The mass of DM is set by considering that it accounts for the entire relic abundance of DM i.e $\Omega_{\text{DM}}h^{2}=0.12$ Aghanim and others (2020) and is found to be $M_{N_{1}}\simeq 3.8$ MeV which is larger than the lower mass bound on feebly interacting massive particles (FIMP) mass which is calculated to be $\simeq 3.5$ MeV according to Eq. 13 as explicitly mentioned in Table 4. To explore the parameter region for successful cogenesis while remaining consistent with the structure formation bounds, we perform an extensive numerical scan utilizing a dedicated Markhov Chain Monte Carlo algorithm following Sarazin et al. (2021) over the 18 free Yukawa couplings and two quartic couplings as given in Table 2. The ranges of the quartic couplings and the Yukawas lie within the theoretical upper bounds obtained from tree-level unitarity Kanemura et al. (1999); Akeroyd et al. (2000); Hambye et al. (2009); Arhrib et al. (2012) and are given by $|\lambda_{j}|\leq 8\pi,Y_{\alpha i}\leq\sqrt{4\pi}\,$, where $\lambda_{j}\in\left\{\lambda_{3},\lambda_{5}\right\}$ and $\alpha,i=1,2$ and 3. We fix the mass scales as $M_{3}=5\times 10^{10}$ GeV, $M_{2}=5\times 10^{9}$ GeV, $M_{\eta}=5\times 10^{8}$ GeV along with the quartic coupling $\lambda_{4}=0$, while the other allowed parameter regions is shown in Fig. 6, where the benchmark point is denoted by ✖. The present and forthcoming collider experiments are not expected to provide a direct probe for the heavy states within this framework. An additional benchmark point with higher values of $\lambda_{3}$ is presented in Appendix A illustrating the dynamics of even larger thermal scattering rate compared to the decay of the inert scalar doublet yielding higher DM mass while remaining safe from the structure formation bounds.